# Invariant derivations and trace bounds

### v3

@@ 18. Minimum values of characters: general case
Now consider b. The identity component $G^{\circ}$ is a normal subgroup of $G$, so its fixed subspace is an invariant submodule. It follows that $G^{\circ}$ either acts trivially or without fixed points. If $G^{\circ}$ acts trivially and the action of $G$ factors through the finite group $G/G^{\circ}$, then the quotient is not cyclic (because $\chi$ is irreducible and $\chi(1)>1$), so b holds by Burnside [Bur03, p. 115]. If $G^{\circ}$ acts without fixed points, then the average value of $\operatorname{Re}\chi$ on $G^{\circ}$ is zero; since $\operatorname{Re}\chi(G^{\circ})$ is connected and takes a positive value $\chi(1)$, it must also take a negative value, and we conclude b. ∎
@@
for $\Phi^{+}$ the positive roots and $s,t$ in the Cartan algebra $\mathfrak{t}(\mathbb{C})$ of the complex form of $G$. By convention, we take the split form, so that the Lie algebra of $T$ is $i{\mathfrak{t}}(\mathbb{R})$. In the right side of (18.4), the products $r\cdot s$ and $r\cdot t$ are canonically determined by the root system and for the inner product $s\cdot wt$ we use the Killing form. The Killing form is nondegenerate [Bou05, §IX.1.3, Prop. 1], so defines an isomorphism of $\mathfrak{t}(\mathbb{C})$ with its dual, and therefore also an inner product among elements of the dual which we use to evaluate $r\cdot\rho$. Up to scaling one of the arguments by $i$, $X$ is the Fourier transform of the Duistermaat-Heckman measure [ER24].
@@
To establish the desired bound on $\min\operatorname{Re}\chi/\chi(1)$, we will show after some work that $X(s,it)$ takes negative values for any nonzero $s\in\mathfrak{t}(\mathbb{R})$ (Lemma 18.13). We write $\lambda$ for the Haar measure on $\mathfrak{t}(\mathbb{R})$, since it is essentially just Lebesgue measure. To avoid issues with the normalization of the Haar measure, the root system, and the Killing form, define a scaling constant
@@ Appendix A Application to finite groups and fields of prime characteristic
Serre’s bounds on the trace of elements of a compact real Lie group can also be used to give bounds on the values of the Brauer character for some representations of finite groups over fields of prime characteristic. Here is one such result:
@@
Let $G$ be an adjoint simple algebraic group over a finite field $k$ of characteristic $p\neq 2$. Then the Brauer character on $G(k)$ is real-valued and its minimum value is $\underline{m}(1)$ as given by Theorem 2.3, unless perhaps $p=3$ and $G$ is of type $A_{2}$ or $E_{6}$.
Here is how we define the Brauer character, compare [Ser77, §18.1] or [Ben91, §5.3]. Let $\Gamma$ be a finite group and $k$ a finite field of characteristic $p\neq 0$. Write the exponent of $\Gamma$ as $\gamma p^{r}$ for $\gamma$ not divisible by $p$. Pick a ring $\hat{\mathcal{O}}\subset\mathbb{C}$ containing the primitive root of unity $\exp(2\pi i/\gamma)$ and with a surjection $\hat{\mathcal{O}}\to\hat{k}$ where $\hat{k}$ is an algebraic extension of $k$. The surjection gives an isomorphism between the group of $\gamma$ roots of unity in $\mathbb{C}$ and those in $\hat{k}$ (because $p$ does not divide $\gamma$), and for each such root of unity $\zeta\in\hat{k}$, we write $\tilde{\zeta}$ for the corresponding root of unity in $\mathbb{C}$.
@@
In the special case where $G$ is a torus, the map $G(\hat{\mathcal{O}})\to G(\hat{k})$ gives an isomorphism on the subgroups of elements of order dividing $\gamma$. For such a $g\in G(\mathcal{O})$, the Brauer character $b_{\rho_{k}}(g)$ equals $\operatorname{Tr}\rho_{\mathbb{C}}(g)$.
@@
Suppose $G$ is a split connected reductive group over $\mathcal{O}$ and put $X$ for a maximal compact subgroup of $G(\mathbb{C})$. Then
$\min\{\operatorname{Re}b_{\rho_{k}}(g)\mid\text{$p$-regular $g\in G(k)$})\}\geq\min\{\operatorname{Re}\operatorname{Tr}\rho_{\mathbb{C}}(g)\mid g\in X\}.$
If $\operatorname{Tr}((\rho_{\mathbb{C}})|_{X})$ is real-valued, then so is $b_{\rho_{k}}$.
@@
Suppose $g\in G(k)$ is $p$-regular. We aim to show that the image of $g\in G(\hat{k})$ is the image of an element $\hat{g}\in G(\hat{\mathcal{O}})$ of the same order. Then $b_{\rho_{k}}(g)=\operatorname{Tr}\rho_{\mathbb{C}}(\hat{g})$. Since elements of finite order in $G(\mathbb{C})$ are conjugate to elements of $X$, this will suffice to prove the claim.
@@
For every representation of $G\times k$, the image of $g$ has minimal polynomial dividing $x^{p}-1$ and in particular the minimal polynomial has distinct roots, so $g$ is semisimple. It follows that $g$ is contained in a maximal torus $T^{\prime}$ of $G$ defined over $k$. Enlarging $\hat{\mathcal{O}}$ and $\hat{k}$ if necessary, we may assume that $\hat{k}$ splits $T^{\prime}$. Therefore, $T^{\prime}\times\hat{k}$ is conjugate under $G(\hat{k})$ to $T\times\hat{k}$. The value of $b_{\rho_{k}}(g)$ only depends on the conjugacy class of $g$ in $G(\hat{k})$, so we may replace $g\in G(\hat{k})$ by its conjugate lying in $T(\hat{k})$ and the proof is complete. ∎
@@
(A.3) $\inf\{\operatorname{Re}b_{\rho_{\hat{k}}}(g)\mid\text{$p$-regular $g\in G(\hat{k})$})\}=\inf\{\operatorname{Re}\operatorname{Tr}\rho_{\mathbb{C}}(g)\mid g\in X\}.$
To see this, note that the elements of $X$ of finite order not divisible by $p$ are dense in $X$. So it suffices to show that each such element is conjugate to an element coming from $G(\hat{\mathcal{O}})$, which follows by the same sorts of argument as in the proof of the lemma.
The “inf” on the right side of (A.3) can be replaced by “min” if and only if the minimum is obtained at an element of $X$ of finite order not divisible by $p$.
@@
One approach to this is via the Weyl group. Put $\mathsf{G}$ for the simply connected cover of $G$ over $\mathcal{O}$. Since $p\neq 2$, 2 is invertible in $\mathcal{O}$ and there is a finite subgroup $W^{*}\subseteq N_{G}(T)(\mathcal{O})$ that is an extension of the Weyl group (as a finite group) by $(\mathbb{Z}/2)^{\operatorname{rank}G}$ and a section $\nu\colon W\to W^{*}$, see for example [Tit66], [Bou05, §IX.4, Exercise 12], or [MPS88]. If $-1$ is in the Weyl group, take $w=-1$. If $G$ has type $D_{n}$ for odd $n\geq 5$ or type $E_{6}$, then we take $w\in W$ to be the element of order 4 or 3 respectively from the proof of Lemma 15.4. In each of these cases, $w$ fixes no root, $\nu(w)$ has order not divisible by $p$, and $\operatorname{Tr}(w|_{\mathfrak{t}})=\underline{m}(1)$. Thus $\nu(w)\in G(\mathcal{O})$ maps to a $p$-regular element of $G(k)$ with Brauer character $\underline{m}(1)$.
@@
Non-split case: Suppose now that $G$ is not split. Since $k$ is finite, $G\times k$ is quasi-split [Mil17, Prop. 17.99], has type ${}^{2}A_{n}$ for $n\geq 2$, ${}^{2}D_{n}$ for $n\geq 4$, ${}^{3}D_{4}$, or ${}^{2}E_{6}$, and is split by an extension $K$ of $k$ of degree 2, 2, 3, or 2. Phrased differently, $G\times k$ is obtained by twisting the split adjoint group $G_{0}$ over $k$ of the same type as $G$ by a semilinear automorphism we denote by $\sigma$ that combines a graph automorphism of order $[K:k]$ of $G_{0}$ with a non-trivial $k$-algebra automorphism of $K$. Since the minimum value of the Brauer character on $G(k)$ is no smaller than that on $G(K)$, it suffices to verify that an element of $G(K)$ whose Brauer character equals $\underline{m}(1)$ belongs to $G(k)$, i.e., is stable under the Galois action.
@@
For $G$ of type $E_{6}$, let $x$ be the Weyl group element from the proof of Lemma 15.4. Since the conjugacy class of $x$ is the unique one of that size, there is a $y\in W$ such that $\sigma(x)=yxy^{-1}$. Put $g$, $h$ for the images of $\nu(x)$, $\nu(y)$ in $G(k)$. Since $g$ normalizes $T$, so does $\sigma(g)$, which acts as $yxy^{-1}$ on $T$ and preserves no roots, so $\sigma(g)=hgh^{-1}$. By Lang’s Theorem [Mil17, Cor. 17.97], there is an element $h_{0}\in G(k^{\mathrm{a}})$ for $k^{\mathrm{a}}$ the algebraic closure of $\hat{k}$, such that $h=\sigma(h_{0})h_{0}^{-1}$. It follows that $h_{0}^{-1}gh_{0}$ is fixed by $\sigma$, i.e., it belongs to $G(k)$. Since it acts as $x$ on $T$, this case is complete.
In the remaining case is where $G$ has type ${}^{2}D_{n}$ for odd $n\geq 5$, we imitate the construction in the split case but noting that the 4-cycle in the Weyl group of $\operatorname{SU}(4)$ exists as described a few paragraphs above. ∎

### v4

@@ 18. Minimum values of characters: general case
Now consider b. The identity component $G^{\circ}$ is a normal subgroup of $G$, so its fixed subspace is an invariant submodule. It follows that $G^{\circ}$ either acts trivially or without fixed points. If $G^{\circ}$ acts trivially, then the action of $G$ factors through the finite group $G/G^{\circ}$, which is not cyclic (because $\chi$ is irreducible and $\chi(1)>1$), so b holds by Burnside [Bur03, p. 115]. If $G^{\circ}$ acts without fixed points, then the average value of $\operatorname{Re}\chi$ on $G^{\circ}$ is zero; since $\operatorname{Re}\chi(G^{\circ})$ is connected and takes a positive value $\chi(1)$, it must also take a negative value, and we conclude b. ∎
@@
for $\Phi^{+}$ the positive roots and $s,t$ in the Cartan algebra $\mathfrak{t}(\mathbb{C})$ of the complex form of $G$. By convention, we take the split form, so that the Lie algebra of $T$ is $i{\mathfrak{t}}(\mathbb{R})$. In the right side of (18.4), the products $r\cdot s$ and $r\cdot t$ are canonically determined by the root system. Since $G$ is compact, there is a $G$-invariant inner product on $\mathfrak{g}$ [Bou05, §IX.1.3, Prop. 1]; pick one and use it to calculate $s\cdot wt$. The inner product defines an isomorphism of $\mathfrak{t}(\mathbb{C})$ with its dual, and therefore also an inner product among elements of the dual which we use to evaluate $r\cdot\rho$. Up to scaling one of the arguments by $i$, $X$ is the Fourier transform of the Duistermaat-Heckman measure [ER24].
@@
To establish the desired bound on $\min\operatorname{Re}\chi/\chi(1)$, we will show after some work that $X(s,it)$ takes negative values for any nonzero $s\in\mathfrak{t}(\mathbb{R})$ (Lemma 18.13). We write $\lambda$ for the Haar measure on $\mathfrak{t}(\mathbb{R})$, since it is essentially just Lebesgue measure. To avoid issues with the normalization of the Haar measure, the root system, and the $G$-invariant inner product, define a scaling constant
@@ Appendix A Application to finite groups and fields of prime characteristic
Serre’s bounds on the trace of elements of a compact Lie group can also be used to give bounds on the values of the Brauer character for some representations of finite groups over fields of prime characteristic. Here is one such result:
@@
Let $G$ be an adjoint simple algebraic group over a finite field $k$ of characteristic $p\neq 2$. Then the Brauer character of $G(k)$ acting on $\operatorname{Lie}(G)$ is real-valued and its minimum value is $\underline{m}(1)$ as given by Theorem 2.3, unless perhaps $p=3$ and $G$ is of type $A_{2}$ or $E_{6}$.
Here is how we define the Brauer character, compare [Ser77, §18.1] or [Ben91, §5.3]. Let $\Gamma$ be a finite group and $k$ a finite field of characteristic $p\neq 0$. Write the exponent of $\Gamma$ as $\gamma p^{r}$ for $\gamma$ not divisible by $p$. Pick a ring $\hat{\mathcal{O}}\subset\mathbb{C}$ containing the primitive root of unity $\exp(2\pi i/\gamma)$ and with a surjection $\hat{\mathcal{O}}\to\hat{k}$ where $\hat{k}$ is an algebraic extension of $k$. The surjection gives an isomorphism between the group of $\gamma$-th roots of unity in $\mathbb{C}$ and those in $\hat{k}$ (because $p$ does not divide $\gamma$), and for each such root of unity $\zeta\in\hat{k}$, we write $\tilde{\zeta}$ for the corresponding root of unity in $\mathbb{C}$.
@@
In the special case where $G$ is a split torus, the map $G(\hat{\mathcal{O}})\to G(\hat{k})$ gives an isomorphism on the subgroups of elements of order dividing $\gamma$. For such a $g\in G(\mathcal{O})$, the Brauer character $b_{\rho_{k}}(g)$ equals $\operatorname{Tr}\rho_{\mathbb{C}}(g)$.
@@
Suppose $G$ is a split connected reductive group over $\mathcal{O}$ and put $K$ for a maximal compact subgroup of $G(\mathbb{C})$. Then
$\min\{\operatorname{Re}b_{\rho_{k}}(g)\mid\text{$p$-regular $g\in G(k)$})\}\geq\min\{\operatorname{Re}\operatorname{Tr}\rho_{\mathbb{C}}(g)\mid g\in K\}.$
If $\operatorname{Tr}((\rho_{\mathbb{C}})|_{K})$ is real-valued, then so is $b_{\rho_{k}}$.
@@
Suppose $g\in G(k)$ is $p$-regular. We aim to show that the image of $g\in G(\hat{k})$ is the image of an element $\hat{g}\in G(\hat{\mathcal{O}})$ of the same order. Then $b_{\rho_{k}}(g)=\operatorname{Tr}\rho_{\mathbb{C}}(\hat{g})$. Since elements of finite order in $G(\mathbb{C})$ are conjugate to elements of $K$, this will suffice to prove the claim.
@@
For every representation of the base change $G\times k$ of $G$ to $k$, the image of $g$ has minimal polynomial dividing $x^{\gamma}-1$ and in particular the minimal polynomial has distinct roots, so $g$ is semisimple. It follows that $g$ is contained in a maximal torus $T^{\prime}$ of $G$ defined over $k$. Enlarging $\hat{\mathcal{O}}$ and $\hat{k}$ if necessary, we may assume that $\hat{k}$ splits $T^{\prime}$. Therefore, $T^{\prime}\times\hat{k}$ is conjugate under $G(\hat{k})$ to $T\times\hat{k}$. The value of $b_{\rho_{k}}(g)$ only depends on the conjugacy class of $g$ in $G(\hat{k})$, so we may replace $g\in G(\hat{k})$ by its conjugate lying in $T(\hat{k})$ and the proof is complete. ∎
@@
(A.3) $\inf\{\operatorname{Re}b_{\rho_{\hat{k}}}(g)\mid\text{$p$-regular $g\in G(\hat{k})$})\}=\inf\{\operatorname{Re}\operatorname{Tr}\rho_{\mathbb{C}}(g)\mid g\in K\}.$
To see this, note that the elements of $K$ of finite order not divisible by $p$ are dense in $K$. So it suffices to show that each such element is conjugate to an element coming from $G(\hat{\mathcal{O}})$, which follows by the same sorts of argument as in the proof of the lemma.
The “inf” on the right side of (A.3) can be replaced by “min” if and only if the minimum is obtained at an element of $K$ of finite order not divisible by $p$.
@@
One approach to this is via the Weyl group. Since 2 is invertible in $\mathcal{O}$, there is a finite subgroup $W^{*}\subseteq N_{G}(T)(\mathcal{O})$ that is an extension of the Weyl group (as a finite group) by a product of copies of $\mathbb{Z}/2$ and there is a section $\nu\colon W\to W^{*}$, see for example [Tit66], [Bou05, §IX.4, Exercise 12], or [MPS88]. If $-1$ is in the Weyl group, take $w=-1$. If $G$ has type $D_{n}$ for odd $n\geq 5$ or type $E_{6}$, then we take $w\in W$ to be the element of order 4 or 3 respectively from the proof of Lemma 15.5. In each of these cases, $w$ fixes no root, $\nu(w)$ has order not divisible by $p$, and $\operatorname{Tr}(w|_{\mathfrak{t}})=\underline{m}(1)$. Thus $\nu(w)\in G(\mathcal{O})$ maps to a $p$-regular element of $G(k)$ with Brauer character $\underline{m}(1)$.
@@
Non-split case: Suppose now that $G$ is not split. Since $k$ is finite, $G$ is quasi-split [Mil17, Prop. 17.99]; has type ${}^{2}A_{n}$ for $n\geq 2$, ${}^{2}D_{n}$ for $n\geq 4$, ${}^{3}D_{4}$, or ${}^{2}E_{6}$; and is split by an extension $K$ of $k$ of degree 2, 2, 3, or 2 respectively. Phrased differently, $G$ is obtained by twisting the split adjoint group $G_{0}$ over $k$ of the same type as $G$ by a semilinear automorphism we denote by $\sigma$ that combines a graph automorphism of order $[K:k]$ of $G_{0}$ with a non-trivial $k$-algebra automorphism of $K$. Since the minimum value of the Brauer character on $G(k)$ is no smaller than that on $G(K)$, it suffices to verify that an element of $G(K)$ whose Brauer character equals $\underline{m}(1)$ belongs to $G(k)$, i.e., is stable under the Galois action.
@@
For $G$ of type $E_{6}$, let $x$ be the Weyl group element from the proof of Lemma 15.5. Since the conjugacy class of $x$ is the unique one of that size, there is a $y\in W$ such that $\sigma(x)=yxy^{-1}$. Put $g$, $h$ for the images of $\nu(x)$, $\nu(y)$ in $G(k)$. Since $g$ normalizes $T$, so does $\sigma(g)$, which acts as $yxy^{-1}$ on $T$ and preserves no roots, so $\sigma(g)=hgh^{-1}$. By Lang’s Theorem [Mil17, Cor. 17.97], there is an element $h_{0}\in G(\bar{k})$ for $\bar{k}$ the algebraic closure of $\hat{k}$, such that $h=\sigma(h_{0})h_{0}^{-1}$. It follows that $h_{0}^{-1}gh_{0}$ is fixed by $\sigma$, i.e., it belongs to $G(k)$. Since it acts as $x$ on $T$, this case is complete.
In the remaining case where $G$ has type ${}^{2}D_{n}$ for odd $n\geq 5$, we imitate the construction in the split case but noting that the 4-cycle in the Weyl group of $\operatorname{SU}(4)$ exists as described three paragraphs above. ∎